Circular No. 7861

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bs IN IC 4221 J. Y. Wei, Y. L. Qiu, H. L. Cao, J. Wang, and J. Y. Hu, Beijing

Astronomical Observatory (BAO), report the discovery of an apparent supernova on unfiltered images taken with the BAO 0.6-m telescope on Mar. 11.73 (mag ∼ 15.5) and 23.73 UT (mag ∼ 14.1). SN 2002bs is located at α = 13^{h}18^{m}30^{s}.30, δ = −14^{o}36′31′′.8 (equinox 2000.0), which is at the center of IC 4221. The unfiltered image taken on Feb. 8 showed no star at the position of the supernova. A low-dispersion spectrum taken with the BAO 2.16-m telescope on Mar. 26.67 shows that SN 2002bs is a type-Ia supernova at about maximum brightness (V ∼ 14.1 from the flux-calibrated spectrum).

XTE J1908+094 M. Feroci, Istituto di Astrofisica Spaziale e Fisica Cosmica, CNR,

Rome; and L. Reboa, BeppoSAX Science Operation Center, Telespazio, Rome, on behalf of the BeppoSAX team, report: “As a part of the followup program on soft γ-ray repeaters, BeppoSAX serendipitously observed the new x-ray transient XTE J1908+094 (IAUC 7856). Due to technical problems with the battery system of the spacecraft, the source was observed only with the Phoswich Detection System (PDS) in the nominal energy range 15–300 keV. The observation started on Mar. 9.386 UT and lasted until Mar. 12.210, for a net observing time of ∼ 48 000 s. Attributing all the counts detected in the collimated field-of-view of the PDS (1^{o}.3, full width at half maximum) to the new source, we can derive a high-energy spectrum for the new x-ray transient. The source is well detected up to > 250 keV, with a net count rate of (18.24 ± 0.04) counts/s at 15–300 keV. The hard x-ray spectrum is not consistent with an extension of the absorbed power law (photon index 1.55) reported for the energy range 2–30 keV (IAUC 7856). Instead, our data require the introduction of a highenergy cut-off at ∼ 100 keV (but still resulting in an unacceptable fit, with χ^{2} of ∼ 5 per degree of freedom). Based on this model, taking into account the 24′ off-axis location of the source, we derive a source flux of ∼ 2.9 × 10^{−9} erg cm^{−2} s^{−1} in the energy range 15–100 keV. The BeppoSAX/PDS detection of the source up to 250 keV strongly supports the interpretation of the new x-ray transient as a new blackhole candidate.”

2002 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green